Pulled it out again today to take another stab, as I slowly work my way through the bin of things that came close to working but not close enough to box up.
Yup, bi-fet quads don't work at all. I used my last 3403 for an OD-1 clone so I can't reproduce what it did, but a second stab at using different Bi-Fets was unsuccessful. Put the 324 back in the socket, and it worked again but has the same issue, namely that there is a very pronounced thunk at the onset of each sweep. Additionally, if the resonance is set higher, it will "lock up" and emit a high-pitched oscillation that won't stop until I turn the sensitivity/Range control was turn way down. Perhaps that can provide some clues as to what is causing the misbehaviour. Could the 220k/100k pair providing a bias/offset to the control current be a possible solution?